In 2.0 minutes, a ski lift raises 4 skiers at constant speed to a height of 135 m. The average mass of each skier is 65 kg. What is the average power provided by the tension in the cable pulling the lift?
Answer #1

here,

number of skiers , n = 4

time taken , t = 2 min = 120 s

mass of each skier , m = 65 kg

height , h= 135 m

the average power provided by the tension in the cable pulling the lift , P = n * m * g * h /t

P = 4 * 65 * 9.81 * 135 /120 W

P = 2.87 * 10^3 W